# PAINT: Parallel-in-time Neural Twins for Dynamical System Reconstruction

### v1

@@ A.2 Implementation and architecture
FlowPAINT and the autoregressive UNet were both implemented in Pytorch [57]. For FlowPAINT we trained in float16 mixed precision. The most important training parameters are provided in Table 3. Our code also uses FlashAttention [17, 16]. For both models we use 20 denoising steps for generating samples.
@@
For the UNet we followed [38] and took the implementation of the public Github repository  [^2] https://github.com/tum-pbs/autoreg-pde-diffusion. We matched the number of parameters by increasing the default model dimension from 128 to 224. For probe-conditioning, we followed the same recipe as for prior timestep conditioning and added mask and probe values as additional channels.
Compute requirements.
The autoregressive baseline is comparably cheap. It was trained on a single Nvidia H100-SXM-80GB for ca. 24 hours. The inference times are reported in the main paper.
@@
History length 16 2
Forward prediction 8 1
@@
Batch size 144 144
@@
Appendix B Theory
B.1 Parallel-in-time models are on-trajectory
Here we provide a simple derivation to show that parallel-in-time models are on-trajectory under Assumption 1. We will show that for an appropriate window size $h$, for all $t\in\mathbb{N}$ and any $\epsilon_{1}>0$:
First, from Assumption 1 choose a window size $h$ such that:
Then, from the Universal Function Approximator Theorem [14], choose an approximation of the true distribution as follows:
B.2 In general, autoregressive models are not on-trajectory
do not generally fulfill the on-trajectory property. We construct this example while demonstrating the over-reliance on autoregressive state. Being on-trajectory is defined as: for all $t\in\mathbb{N}$ and any $\epsilon>0$:
From Section 3.3 it is known that autoregressive models accumulate errors with the product of Jacobians, similar to classical ODEs [55].
Logistic map.
Let $f:\mathbb{R}\mapsto\mathbb{R}$ be the logistic map, which is chaotic for $r=3.8$ [52]:
$f(x_{t-1},\eta_{t})$ $=r\cdot x_{t-1}\cdot(1-x_{t-1})$ (13)
Appendix C Additional results
We show additional results for other Reynolds numbers and probe constellations in Figures 9, 10, 11, 12, 14, 13, 15, and 16.
Figure 9: Left: Exemplary groundtruth velocity snapshot, for Reynolds number 2100 and using the vertical probe point constellation. Center: Mean of the predictions over 10 independent seeds. Right: Standard deviation of the predictions over the same 10 seeds. All values are normalized by the mean inlet velocity.
@@ Appendix C Additional results
Figure 10: Statistical analysis of reconstructed flow fields for Reynolds number 2100 and a grid probe point constellation.
@@
Figure 11: Statistical analysis of reconstructed flow fields for Reynolds number 1100 and a grid probe point constellation.
@@
Figure 12: Statistical analysis of reconstructed flow fields for Reynolds number 1100 and a vertical probe point constellation.
(a) Vertical probe constellation.
(b) Grid probe constellation.
Figure 13: Histogram of all MSE values of the Reynolds number 2100 test trajectory and a grid probe point constellation.
(a) Temporal evolution of the mean and standard deviation of the MSE when comparing the parallel-in-time and autoregressive approaches.
(b) Histogram of all MSE values accumulated over the rollout, comparing the parallel-in-time and autoregressive approaches.
Figure 14: Comparison of error characteristics between the parallel-in-time and autoregressive reconstruction approaches for Reynolds number 1100 and a grid probe point constellation.
(a) Temporal evolution of the mean and standard deviation of the MSE when comparing the parallel-in-time and autoregressive approaches.
(b) Histogram of all MSE values accumulated over the rollout, comparing the parallel-in-time and autoregressive approaches.
Figure 15: Comparison of error characteristics between the parallel-in-time and autoregressive reconstruction approaches for Reynolds number 1100 and a vertical probe point constellation.
@@
Figure 16: The top rows show five exemplary LES ground truth snapshots. Time steps with $t\leqslant 0$ correspond to reconstructions based on measurements, while time steps with $t>0$ represent predictions with no measurements present. The following three rows display reconstructions from three independent random seeds, each showing a connected sequence generated based on the probe information provided. All values are normalized by the mean inlet velocity.

### v2

@@ B.2 Implementation and architecture
FlowPAINT and the autoregressive UNet were both implemented in Pytorch (Paszke et al., 2017). For FlowPAINT we trained in float16 mixed precision. The most important training parameters are provided in Table 3. Our code also uses FlashAttention (Dao et al., 2022; Dao, 2024). For both models we use 20 denoising steps for generating samples.
@@
For the UNet we followed (Kohl et al., 2023) and took the implementation of the public Github repository  [^2] https://github.com/tum-pbs/autoreg-pde-diffusion. We matched the number of parameters by increasing the default model dimension from 128 to 224. For probe-conditioning, we followed the same recipe as for prior timestep conditioning and added mask and probe values as additional channels. We trained two UNet variants, one conditioned on probe measurements from the current timestep and another conditioned on a 16-timestep probe history, to ensure a fair comparison with FlowPAINT as well as demonstrate that even an extensive probe history cannot overcome the dominance of the autoregressive state in a given prediction.
Autoregressive Diffusion Transformer.
We tested the Diffusion Transformer (Peebles & Xie, 2023) architecture used for PAINT as an autoregressive model. However, we observe that its predictive performance degrades rapidly after only a few rollout steps (see Figure 22). Consequently, we rely on the above-mentioned well-established UNet architecture mentioned above for the primary comparison to ensure a fair and robust baseline.
Optimization hyperparameters.
We employed well-established techniques and hyperparameters for optimization, including AdamW (Loshchilov et al., 2017), cosine annealing of the learning rate, etc. These choices were grounded in their proven effectiveness and widespread adoption in the literature. Although we acknowledge that tuning these hyperparameters could potentially yield marginal performance improvements, we do not anticipate that such modifications would substantially impact the primary conclusions and insights drawn from our experiments. Given the computational constraints and the dimensionality of the hyperparameter search space, we chose to use the established settings for the sake of experimental tractability. Future work may explore the potential benefits of hyperparameter tuning in this context to validate this assumption.
Selection of batch size and timesteps per batch.
As mentioned in Section 4, a direct comparison between PAINT and the autoregressive baseline is not entirely feasible, given PAINT’s prediction of a full sequence of states. Therefore, we carefully evaluated and chose the batch size and the timesteps per batch according to the following procedure
For the autoregressive UNet, we opted for the maximum batch size that could be accommodated by a single GPU, which amounted to 144. We also tested a batch size of 288 on 2 GPUs via data-parallel training. However, as there was no significant change in training nor validation loss, we settled on a batch size of 144 for all experiments.
In the case of FlowPAINT 16/8, we maintained the same batch size of 144. For FlowPAINT 16/0, we tested two variants: one with a batch size of 144, and another with an increased batch size of $144\times 1.5=216$. The latter was chosen to align the number of predicted timesteps with the 16/8 setting. Given the negligible difference observed between the two variants, we proceeded with a batch size of 144 for the remainder of the FlowPAINT experiments.
@@
Autoregressive history length 0 2
Probe history length 16 [16, 1]
Forward prediction [8, 0] 1
@@
Batch size [144, 216] 144
Compute requirements.
The autoregressive baseline is comparably cheap. It was trained on either a single Nvidia RTX-PRO6000 Blackwell 96GB or a single Nvidia H100-SXM-80GB for ca. 24 hours. The inference times are reported in the main paper.
@@
B.3 Additional results and ablations
We perform ablations to gain further insights on the importance of certain components. Ablation results are summarized in Table 5, with further evaluations shown in Figures 19, 20, 17, 18 and 22.
Stability through future predictions.
Comparing the smoothing-and-filtering variant, FlowPAINT 16/0 with the smoothing-filtering-and-prediction variant, FlowPAINT 16/8 highlights the stabilizing effect of incorporating future-state predictions. While the FlowPAINT 16/0 approach relies exclusively on reconstructing the given measurement series, the inclusion of predictive task seems to enable FlowPAINT 16/8 to produce more stable reconstructions. FlowPAINT 16/0 performs in general worse than FlowPAINT 16/8, though still better than the autoregressive baseline.
Figure 9: The solid lines indicate the mean MSE over a reconstructed sequence, with the shaded regions indicating ± one standard deviation across samples. While the average MSE for future timesteps remains relatively low, FlowPAINT surprisingly exhibits the highest error for timesteps in the past. This data stems from an evaluation using Reynolds number 2100 and the grid probe point constellation.
Figure 10: MSE over number of probe points provided during inference. Probe points are arranged in a pattern comparable to the grid probe constellation. For FlowPAINT 16/8, the MSE decreases as the number of probe points increases, reaching an optimum near 100. Beyond this point, the MSE increases again, as FlowPAINT was only trained with 25 probe points. In contrast, the performance of the autoregressive baseline shows little sensitivity to the number of probe points provided.
Model Cond. Batchsize ReRe Probes ${\text{MAE}}(\overline{\mathbf{u}})\downarrow$ ${\text{MAE}}(\overline{\mathbf{u}^{\prime 2}})\downarrow$ $\text{MSE}([\mathbf{\mathbf{u_{[0,t]}}}])\downarrow$ $\text{RMSE}(E(k))\downarrow$
$\times 10^{-3}$ $\times 10^{-5}$ $\times 10^{-5}$ $\times 10^{-6}$
FlowPAINT 16/8 144 2100 grid $1.2\pm 1.6$ $9.9\pm 21.7$ $8.1\pm 3.6$ $6.9$
16/0 144 $1.3\pm 2.6$ $8.7\pm 15.1$ $10.8\pm 3.9$ $6.3$
16/0 216 $1.5\pm 3.2$ $9.1\pm 15.5$ $11.5\pm 4.9$ $6.9$
Autoreg. 16/0 144 2100 grid $8.2\pm 13.7$ $93.0\pm 103.2$ $155.0\pm 50.1$ $66.3$
1/0 144 $9.5\pm 14.3$ $98.7\pm 110.4$ $158.3\pm 46.8$ $70.9$
FlowPAINT 16/8 144 2100 vertical $2.2\pm 3.1$ $30.2\pm 38.4$ $49.0\pm 21.0$ $20.1$
16/0 144 $3.0\pm 3.2$ $30.0\pm 41.6$ $68.2\pm 25.3$ $20.5$
16/0 216 $2.9\pm 2.9$ $25.7\pm 34.6$ $70.6\pm 26.6$ $18.9$
Autoreg. 16/0 144 2100 vertical $7.4\pm 12.5$ $87.6\pm 96.1$ $156.2\pm 50.9$ $62.4$
1/0 144 $9.2\pm 13.8$ $97.8\pm 108.8$ $157.3\pm 70.0$ $71.8$
FlowPAINT 16/8 144 1100 grid $0.9\pm 1.3$ $2.2\pm 5.5$ $\pm 1.0 2.02.0$
16/0 144 $1.2\pm 2.8$ $6.2\pm 10.7$ $\pm 1.6 4.64.6$
16/0 216 $1.0\pm 2.2$ $4.9\pm 8.2$ $3.8\pm 1.3$ $3.4$
Autoreg. 16/0 144 1100 grid $3.7\pm 6.3$ $16.0\pm 24.1$ $29.9\pm 8.8$ $12.3$
1/0 144 $3.5\pm 5.6$ $17.3\pm 25.9$ $28.8\pm 8.3$ $12.4$
FlowPAINT 16/8 144 1100 vertical $1.1\pm 1.6$ $6.3\pm 11.1$ $8.9\pm 3.4$ $3.7$
16/0 144 $3.0\pm 5.8$ $8.2\pm 15.8$ $24.2\pm 6.5$ $9.9$
16/0 216 $3.4\pm 6.2$ $10.2\pm 17.5$ $27.2\pm 6.9$ $11.7$
Autoreg. 16/0 144 1100 vertical $3.7\pm 6.1$ $15.9\pm 23.9$ $29.5\pm 8.8$ $12.3$
1/0 144 $3.4\pm 5.5$ $16.9\pm 25.5$ $28.8\pm 8.1$ $12.2$
(a) Vertical probe constellation.
(b) Grid probe constellation.
Figure 11: MSE over time of ground-truth vs. predicted trajectories for the $Re=1100$ test trajectory and different probe point constellations, comparing the autoregressive baselines with FlowPAINT 16/8 and 16/0. The autoregressive model indicates over-reliance on the autoregressive state.
Figure 12: Left: Exemplary groundtruth velocity snapshot, for Reynolds number 2100 and using the vertical probe point constellation. Center: Mean of the predictions using FlowPAINT 16/8 over 10 independent seeds. Right: Standard deviation of the predictions using FlowPAINT 16/8 over the same 10 seeds. All values are normalized by the mean inlet velocity.
@@ B.3 Additional results and ablations
Figure 13: Statistical analysis of reconstructed flow fields using FlowPAINT 16/8 for Reynolds number 2100 and a grid probe point constellation.
@@
Figure 14: Statistical analysis of reconstructed flow fields using FlowPAINT 16/8 for Reynolds number 1100 and a grid probe point constellation.
@@
Figure 15: Statistical analysis of reconstructed flow fields using FlowPAINT 16/8 for Reynolds number 1100 and a vertical probe point constellation.
(a) Time-averaged kinetic energy spectrum.
(b) Histogram of the kinetic energy.
Figure 16: Statistical analysis of reconstructed flow fields using FlowPAINT 16/8 for Reynolds number 2100 and a vertical probe point constellation.
(b) Histogram of time-averaged kinetic energy.
(c) Time-averaged kinetic energy spectrum.
Figure 17: Statistical analysis of reconstructed flow fields using FlowPAINT 16/0 for Reynolds number 2100 and a grid probe point constellation.
(b) Histogram of time-averaged kinetic energy.
(c) Time-averaged kinetic energy spectrum.
Figure 18: Statistical analysis of reconstructed flow fields using FlowPAINT 16/0 for Reynolds number 2100 and a vertical probe point constellation.
(b) Histogram of time-averaged kinetic energy.
(c) Time-averaged kinetic energy spectrum.
Figure 19: Statistical analysis of reconstructed flow fields using FlowPAINT 16/0 for Reynolds number 1100 and a grid probe point constellation.
(b) Histogram of time-averaged kinetic energy.
(c) Time-averaged kinetic energy spectrum.
Figure 20: Statistical analysis of reconstructed flow fields using FlowPAINT 16/0 for Reynolds number 1100 and a vertical probe point constellation.
@@
Figure 21: The top rows show five exemplary LES ground truth snapshots. Time steps with $t\leq 0$ correspond to reconstructions based on measurements, while time steps with $t>0$ represent predictions with no measurements present. The following three rows display reconstructions and predictions using FlowPAINT 16/8 from three independent random seeds, each showing a connected sequence generated based on the probe information provided. All values are normalized by the mean inlet velocity.
Figure 22: Autoregressive rollout of a baseline model with an architecture comparable to our PAINT implementation. The prediction quality deteriorates almost immediately, motivating the use of a well-established autoregressive baseline to enable a fair comparison between traditional autoregressive methods and our parallel-in-time approach.
Figure 23: Qualitative illustration of the discontinuities induced by our parallel-in-time approach, using Autoreg. 1/0 and FlowPAINT 16/8 for Reynolds number 2100 and the vertical probe point constellation. Shown are consecutive prediction timesteps of a groundtruth trajectory, alongside the autoregressive baseline and the parallel-in-time approach. Due to its inherently temporally discontinuous nature, PAINT exhibits larger frame-to-frame changes. The autoregressive approach, on the other hand, displays frame-to-frame variations comparable to the ground truth, as each prediction is conditioned on previous outputs.